# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v3

@@ VII-A1 Aleatoric methods
DDPM Flexible, expressive Sequential sampling [98, 99, 100, 101, 94, 95, 96, 97]
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification. Nonetheless, the shortcomings of VAEs are well known. For example, such models suffer from inference suboptimality related to ELBO optimization [207, 65] and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse [73, 63, 68]. This has been hypothesized to be caused by excessively strong decoders [64] and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective [74], residual connections [74, 75], or deep supervision [75] are required for generalization. A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions. Hence, VAE-based models serve as good choice for the task given their shortcomings are sufficiently addressed and partially solved.
Compared to the VAE-based models, the adoption of DDPMs is rather limited. This is regardless of the fact that it outperforms the VAE-based methods. Besides the point that DDPMs emerged much later in the field, their crucial limitation, similar to thee PixelCNN [53], is the tedious sequential inference procedure [92, 208, 209]. This shortcoming is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Regardless of these shortcomings, DDPMs are extremely flexible and avoid loss of crucial high-frequency information often found in latent-variable modeling with dimensionality reduction (e.g. blurry reconstructions of VAEs). Furthermore, it can be noted that the best-performing DDPM models are discrete in nature. While it is debatable whether shifting to categorical distributions is required for complex image generation [136], this observation implies the need for further investigation of the merits of categorical distributions in segmentation settings. Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13).
(a) LIDC-IDRI [131]
(b) CityScapes [132].
Figure 13: Continuous vs. categorical forward diffusion process with cosine noise scheduling [210]. Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
Implicit methods such as TTA [103, 104, 105, 106] have received some use in literature, but have been quickly surpassed by alternative methods. However, TTA does lead in the category of simplicity, requiring no additional mechanisms or modifications to the employed architecture. Furthermore, CAR [56] is a lesser used method and this is likely due to well-known training problems of GAN-based models [211]. Due to this, GANs require additional heuristic terms in the training objective due to instability. For example, the training objective of CAR is a summation of four separate losses.
SSNs [50] directly model correlations in pixel space and deem simple, fast and model-agnostic modeling. Notably, SSNs also suffer from training instability due to the invertibility requirement of the covariance matrix. This can be often simply circumvented by masking out the background to avoid exploding variances and by using uncorrelated Gaussians for datapoints where the covariance matrix is singular. In practice, nevertheless, this quick-fix solution is much more frequently required than seems to be suggested, inviting further research on explicitly modeling the likelihood function in pixel space.
VII-A2 Epistemic methods
Almost all discussed literature quantifying epistemic uncertainty employ approximations of VI. In fact, usage of MC Dropout dominates, mainly due it being a relatively simple, inexpensive and straightforward approach. Nevertheless, MC dropout has been subject to substantial criticism [212, 213, 109, 123]. For example, it has been shown that MC dropout can assign a probability of zero to the true posterior, which can also erroneously possess multi-modality [212]. Furthermore, MC Dropout can be heavily reliant on the interaction between model size and dropout rate, rather than the observed data [213, 109]. These weaknesses have been used as a basis for alternative techniques, where the dropout rate is learned [109, 123]. Ultimately, uncertainty from MC dropout and similar methods should be viewed as an added benefit rather than the main focus of an already functional model. If accurate uncertainty quantification is critical, MC dropout should be avoided in any case.
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling [189, 162], while in other works ensembling excels [163]. Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout when evaluating with the PAvPU metric [151]. All things considered, the preference for a particular methodology seems to carry a strong data-dependency [125, 14]. Our recommendation is to experiment with both explicit VI and its approximations.
VII-B Discussion on applications
Up until now, theoretical insights have mostly been discussed. In this section, a deeper dive is taken into the applications of uncertainty quantification. Table III provides an overview of these applications in various domains. Notably, most models are employed in healthcare use cases, where ambiguity frequently arises due to the trade-off between accuracy and the incisiveness of medical diagnosis systems. Furthermore, in the automotive industry, sensors often operate under constraints, and objects of interest are typically at a considerable distance which induces ambiguity in the acquired images. It is evident that in these fields, uncertainty is primarily utilized to quantify observer variability and/or to correlate uncertainty with prediction accuracy. Research on improved generalization and uncertainty-based active learning is much sparser. In the following sections, each respective downstream task is further discussed.
TABLE III: Overview of domains using uncertainty quantification for segmentation.
Observer Model Active Model
variability Introspection Learning generalization
Lung [50, 69, 59, 73, 60] [171, 177, 189] [182] [189]
[62, 68, 74, 75, 138]
[140, 141, 94, 139, 79]
Brain [50, 98, 96, 142] [163, 159, 161, 160, 177] [197] [159, 188]
[99, 139, 144, 94] [161, 177]
Outdoor scenes [74, 50, 147] [152, 153, 157, 151, 181] [202, 204] [151, 181, 152]
[101, 156] [186, 187]
Cardiovascular [165, 166, 176, 163] [194] [214]
[125, 167, 178]
Prostate [145, 76, 75] [163] [194]
Prostate [138, 139]
Eye [100, 99, 60] [168, 169] [168]
Skin [100] [171, 172] [195, 198]
Indoor scenes [101, 156] [152, 157] [202, 205] [152]
Microscopy [97, 101, 74] [180, 175] [175]
Others [141, 73, 99, 59] [172, 173, 177, 19, 174] [194, 200, 96] [190]
[96, 139]
VII-B1 Observer variability
Varying delineation hypotheses present themselves as “noise” in the ground-truth of the data. This observer variability is often a result of viewing limitations and therefore heavily correlate to the input image. In turn, this explains the success of directly learning input-conditional observer variability, through explicitly modeling the likelihood distribution, rather than inferring the underlying latent parameter distribution. Nonetheless, this approach does have its limitations. For example, it has been shown that such models, without explicit conditioning, do not encapsulate more subtle variations such as distinct labeling styles [215].
A significant challenge in this domain is the lack of standardization, making benchmarking extremely difficult. The reasoning for this is twofold. Firstly, it is evident from Table III that a wide range of datasets are used in literature, often involving proprietary in-house data. This inherently makes replicating the presented results impossible. Secondly, consensus on data processing is also lacking on the publicly available datasets. For example, significant incongruity across literature is observed for the LIDC-IDRI dataset, regarding data splitting and preprocessing. Similarly, there is a lack of agreement on the implementation for the GED and HM metric. Several factors, including the kernel choice, the number of predicted samples, and the handling of empty segmentations, can significantly impact the resulting quantitative evaluations.
Regarding the evaluation metrics, we find that the literature often places excessive emphasis on improving them. This essentially forces the algorithm to predict segmentations identical to the available ground-truth masks. This can be problematic, as it defeats the original goal to predict plausible unseen segmentations, becoming a textbook case of Goodhart’s law: ”When a measure becomes a target, it ceases to be a good measure”. Furthermore, the quality of the GED evaluations have been subject to substantial criticism as well [216, 74, 63]. Hence, practitioners should be extra vigilant and consider the importance to inquire domain-level experts for qualitative evaluation. Because the GED and HM are dependent on on the number of segmentation masks available, an alternative is to involve additional annotators per datapoint for more accurate evaluation, but is likely to be equally expensive. This aspect calls for systematic procedures to evaluate uncertainty subject to limited ground-truth masks.
VII-B2 Model introspection
Literature that evaluates models by correlating model uncertainty with error is, comparatively, much more thorough and standardized. This is especially evident by the frequent use of the PAvPU metric. Furthermore, a majority of the available research use MC Dropout and correlate the output entropy to the prediction accuracy. The disadvantage of relying on MC Dropout when the uncertainty is crucial, has been discussed in Section VII-A2. Therefore, we recommend incorporating explicit VI at specific points in the network to achieve more accurate uncertainty quantification. These points can be selected based on the layers that influence the output segmentation most significantly.
VII-B3 Model generalization
Epistemic uncertainty quantification methods can result in improved performance, as it often serves as a surrogate for model combination. Therefore, uncertainty quantification can be readily available in the toolbox of deep learning researchers together with other commonly used regularization techniques. In this case, simple methods such as ensembling and MC dropout, are despite their criticisms relatively harmless. Furthermore, a parallel characterization can be made with MC Dropout and placing an $L_{2}$-norm penalty on the model weights. While this benefit is compelling, improved model performance can be also obtained with more computationally efficient regularizers. Therefore, improved performance should not be the ultimate goal of uncertainty quantification, but considered as an ancillary advantage.
VII-B4 Active Learning
The challenging task of Active Learning holds a significant promise and can potentially reduce the need for intensive labeling procedures, which often require specialized expertise. Furthermore, this kind of approach points towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate the adoption of such models in sectors requiring extensive specialization. Furthermore, Active Learning can accelerate privacy-centric collaboration when combined with a federated setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local models.
However, Active Learning requires a well-generalized model using little data. This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and typically requires datasets of substantial size. Furthermore, many uncertainty-based approaches regularly extend conventional Active Learning for classification to segmentation by simply aggregating the pixel metrics together. For imbalanced settings, it has been shown that it can perform even worse than random selection [217]. Therefore, additional modifications such as target, boundary or diversity awareness [200, 205], or region-based annotation [204, 218] are often required to apply Active Learning to segmentation tasks.
VII-C Future work
In this penultimate section, we will provide recommendations for future work.
VII-C1 Exploring generative models
Generative modeling, a rapidly evolving field, has successfully been employed for quantifying observer variability. The benefits of quickly applying its developments to segmentation problems have been evident with state-of-the-art DDPMs [90], which were initially proposed for unsupervised image generation. Also, literature on unsupervised VAEs have greatly benefited the PU-Net [73, 63, 68]. Therefore, we advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models. In fact, any unsupervised generative model can be theoretically translated to the supervised setting through intricate conditioning and can be therefore used as a probabilistic segmentation model.
Given this flexibility, it remains unclear why Normalizing Flows (NFs) remain underutilized for this task. Specifically continuous NFs, which approximate the time-dependent score function, strongly resemble DDPMs which have been extensively used for segmentation problems. Furthermore, NFs enable explicit and exact evaluation of the likelihood, which can aid in further interpretation of model predictions. Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian, so that inference times are therefore much faster. However, it should be noted that the invertibility constraint greatly hinders the expressiveness of intermediate functions, leading to memory-intensive architectures.
VII-C2 Variational Inference and beyond
@@
A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred over explicit VI). Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section V-C) for language modeling [219]. Future work should also experiment with this.
@@
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace Approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions. To the best of our knowledge, these approaches have not been studied within the context of uncertainty quantification in segmentation.
VII-C3 Single-pass uncertainty
The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards deterministic uncertainty models [220, 221, 222], which only depend on a single forward pass. Mukhoti et al. [220] show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. This approach achieves faster computation, while also providing both epistemic and aleatoric uncertainty.
@@ VII-C3 Single-pass uncertainty
Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) [223]. While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems [224], and Ancha et al. [225] recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model. Unfortunately, further research has not been published beyond this.
VII-C4 Improved CNN architectures
With the introduction of Vision Transformers [226], CNN-based models have been challenged in the task of semantic segmentation [227, 228, 229]. Regardless the success of Vision Transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture [39]. This is mainly because CNNs already possess desirable inductive biases, while Transformers conversely require extensive pretraining with large datasets [230]. Nonetheless, CNNs have benefited from the recent developments in Transformers. For instance, “ConvNext” takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains [231]. Since many innovations in this field focus on the technique of uncertainty quantification, the applied backbones receive less attention and might gradually become outdated. Our recommendation suggests improving current models with developments in generic CNN-based architectures.
VII-C5 Distributing-free modeling
@@
A distribution-free framework known as Conformal Deep Learning, produces prediction sets that guarantee to contain the ground truth with a user-defined probability. With the help of an additional calibration set, a heuristic notion of ambiguity (i.e. miscalibrated softmax outputs) is transformed to rigorous uncertainty. This technique is especially renown for being model-agnostic, simple and highly flexible [232]. Quite recently, Conformal Prediction has been applied to segmentation problems [233, 234, 235, 236], thereby enjoying the aforementioned benefits and showing an increased traction of this framework. We recommend further research in this direction to discover novel applications and benchmark against current architectures.
VII-C6 Volumetric segmentation
In many clinical datasets, the volumetric data are in most instances sliced to patches and processed with conventional 2D-based CNN-models. However, such models are often straightforward extensions of existing 2D models (and almost exclusively VAE-based), rarely addressing novel challenges introduced with additional dimensionalities. For example, 3D extensions of the PU-Net simply remain to use similar techniques for inserting latent samples in the decoding networks. Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization [179]. Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to implement models for volumetric segmentations problems.
Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models. For example, the GED cannot be estimated volumetrically with 2D models. Moreover, translating the already computation-intensive models (e.g. based on DDPMs or PixelCNNs) to 3D is another challenge, due to the memory requirement involved with volumetric data, which will increase the training load and inference time even further. Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underline the need of more investigation related to 3D uncertainty quantification, that more closely aligns with real-world clinical practice and therefore encourages faster adoption.
VIII Conclusion
Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability in their predictions. With the vast body of literature, encompassing diverse applications and modalities, the need for a comprehensive and systematic overview of the field is addressed. We present clear definitions and unified notation of methodologies that attempt uncertainty modeling, by considering the field from a theoretical perspective. Aleatoric uncertainty can be modeled in pixel or latent space with generative models, or implicitly be expressed with test-time augmentation. Epistemic uncertainty is captured with Variational Inference on the parameter distribution, or approximated with Monte Carlo Dropout or model ensembling. Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks, which enable to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and to the real-world applications.
Our future recommendations pertain aligning the field with advancements in general generative modeling and Deep Learning architectures. Furthermore, the adoption of deterministic uncertainty quantification methods that avoid multiple forward passes, such as Conformal and Evidential Deep Learning, are suggested. Particularly the latter approach is interesting due to its ability to incorporate both uncertainty types. Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques, as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation, is a beneficial contribution to the field. Finally, due to the clinical relevance of uncertainty in semantic segmentation, more attention is advised to models catering to volumetric data. In this manner, the review paper guides researchers on the topic of probabilistic segmentation and suggests future endeavors within the lightning-fast evolving field of Deep Learning-based Computer Vision.

### v4

@@ 7.6 Modeling feature distributions
DDPM Flexible, expressive Sequential sampling (Chen et al., 2023; Zbinden et al., 2023; Rahman et al., 2023)
7.7 Modeling parameter distributions
In the previous sections, we reviewed feature modeling methods, typically used to quantify observer variability. Other downstream applications usually rely on epistemic uncertainty modeling, i.e. modeling parameter distributions, which will be discussed in this section.
Criticism of MC Dropout
Nearly all literature on epistemic uncertainty quantification relies on approximations of variational inference (VI). Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation. However, MC Dropout has faced substantial criticism Folgoc et al. (2021); Osband (2016); Kingma et al. (2015); Gal et al. (2017a). For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality Folgoc et al. (2021). It is also sensitive to model size and dropout rate rather than the observed data Osband (2016); Kingma et al. (2015). To address these issues, some alternatives attempt to learn the dropout rate Kingma et al. (2015); Gal et al. (2017a). Still, uncertainty estimates from MC Dropout should be treated as a convenient byproduct—not a reliable core feature. When precise uncertainty quantification is critical, such methods are generally insufficient and best avoided. When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary. In such cases, methods like ensembling and MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation. However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors. For instance, both ensembling and MC Dropout introduce a model combination effect—effectively averaging multiple networks. MC Dropout also resembles placing an $L_{2}$ penalty on weights. While these effects can improve generalization, similar gains might be achieved with simpler, more efficient regularizers. Perhaps it is the stochasticity itself that helps. Still, if performance is the main goal, uncertainty modeling is likely not the most efficient route due to its added computational cost. Instead, improved performance may simply be a side effect of these techniques.
Comparative studies
@@ 7.7 Modeling parameter distributions
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling Hoebel et al. (2020); Roy et al. (2018b), while in other works ensembling excels Mehrtash et al. (2020); Ng et al. (2022). Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout Mukhoti & Gal (2018). All things considered, the preference for a particular methodology seems to carry a strong data-dependency Dahal et al. (2020); Jungo & Reyes (2019). A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred). Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section LABEL:subsec:ME) for language modeling (He, 2024). Future work should also experiment with this.
Variational Inference Alternatives
@@
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace Approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions. Zepf et al. (2023b) explored the impact of the Laplace Approximation on segmentation networks; however, beyond this work, the method has received limited attention in the field.
Sampling-free uncertainty
@@
The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards sampling-free uncertainty models (Mukhoti et al., 2023; Liu et al., 2020; Van Amersfoort et al., 2020; Postels et al., 2019), only depending on a single forward pass. Mukhoti et al. (2023) show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) (Dempster, 1967). While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems (Sensoy et al., 2018), and Ancha et al. (2024) recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model. However, follow-up research in this direction remains limited to date.
Reliable uncertainty estimates
@@
A distribution-free framework known as Conformal Prediction (CP) produces prediction sets that guarantee inclusion of the ground truth with a user-specified probability. By using an additional calibration set, CP transforms heuristic model confidence scores into rigorous uncertainty estimates. This technique is particularly renowned for being model-agnostic, simple, and highly flexible (Angelopoulos & Bates, 2021). Recently, CP has been applied to segmentation tasks (Wieslander et al., 2020; Brunekreef et al., 2024; Mossina et al., 2024; Wundram et al., 2024a), benefiting from these advantages and reflecting the growing interest in the framework for structured prediction. We encourage further research to explore novel applications and benchmark CP against existing architectures.
8 Conclusion
Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability of their predictions. Given the vast and diverse body of related literature spanning various applications and modalities, often scattered and marked by inconsistent evaluation and benchmarking practices, there is a clear need for a comprehensive and systematic overview. This work addresses that need by presenting clear definitions and a unified notation for methodologies in uncertainty modeling. Furthermore, confusion and misunderstanding around uncertainty modeling and quantification have led to diverging beliefs within the field. By structuring the methodologies by targeted uncertainty type and followed by downstream pipeline into four distinct tasks, it becomes easier for researchers to navigate the field and build upon existing work. We place particular emphasis on addressing key conceptual gaps to support both pragmatic and theoretically sound research. We also highlight the potential and broader impact of advancing uncertainty-based active learning. A noticeable trend is the adaptation of generative modeling advances to segmentation through novel, task-specific conditioning strategies for modeling ambiguity in segmentation masks. While such efforts are already underway, it remains important to closely follow progress in general generative machine learning and continue exploring its effective application to segmentation. Finally, we recommend future work to focus on improving spatial aggregation methods, standardization in benchmarking, backbone architectures, as well as the adoption of novel uncertainty estimation techniques, as many commonly used methods lack rigor. In particular, there is growing interest in distribution- and sampling-free approaches. In this manner, this review paper aims to guide researchers toward building more reliable, efficient, and uncertainty-aware segmentation models within the lightning-fast evolving field of Deep Learning-based Computer Vision.